# Meta-Reasoning Improves Tool Use in Large Language Models

### v1

@@ A.3 Inference-time Hyperparameters
Figure 3: Proportion of decoded golden tools across the FuncQA validation set, for k=1, k=5 and k=10, where k is the number of top-probability tokens decoded at each decoding step.
Appendix B Tool Sampling Experiments
Temperature Sampling  We experiment with temperature sampling and find that it is not an optimal strategy to gather candidate tools, as lower temperatures do not lead to enough diversity while higher ones generate irrelevant tools.
Top-k Decoding  We run ToolkenGPT (in its original implementation based on the first version of LLaMA) on the validation set of FuncQA. We find that, in samples where the system has generated an incorrect tool, the correct one – complete with correct arguments – is among the top five most likely tokens in over $60\%$ of cases (we use greedy decoding as in the original paper). These cases include samples where the correct tool can be found by searching the top tokens at a different position from the one that has produced the incorrect tool. Overall, the system decodes the correct tool (regardless of the arguments it generates for it) in 76.9% of samples; this rises by over $10\%$ to $87.2\%$ when we consider the top $k=5$ tokens at each decoding step (Fig. 3). Further increasing $k$ to $10$ incurs a higher computational cost without raising the proportion of decoded golden tools.
@@ Appendix C Baselines Implementation
All our baselines use Llama 3 8B Instruct as the base model.
Llama 3 8B Instruct  We measure our base model’s performance in the zero-shot setting. We experiment with CoT zero-shot prompting (prepending to the question the instruction Let’s think step by step) but find that just using the raw question as input results in better performance.
Chain-of-Thought (CoT)  We extract from the training set six pairs of questions and answers demonstrating maths operations and their use. For comparability, we use the same exemplars as in the implementation of ToolkenGPT and Tecton. Note that Llama 3 8B Instruct achieves $79.6\%$ accuracy Dubey et al. (2024) on the non-enhanced version of GSM8K when prompted in few-shot CoT fashion. Our experiments show that it performs significantly worse ($37.3\%$) on GSM8K-XL, highlighting the difficulty of the enhanced dataset.
Trice  We train Trice using the same hyperparameters as in the original implementation, with the exception of the batch size in the first phase of training, reduced to 128 with eight gradient accumulation steps due to the memory limitations of our hardware. Since we train on a single dataset (GSM8K-XL), our implementation corresponds to the setup referred to as Trice-split in the original paper. Note that Qiao et al. (2024) implement Trice on top of Alpaca Taori et al. (2023), ChatGLM Zeng et al. (2024) and Vicuna Zheng et al. (2023). For comparability with Tecton and the other baselines, here we use Llama 3 as the base model. We thus adjust Trice’s prompt templates and special tokens to be consistent with Llama 3’s model card [^1] https://www.llama.com/docs/model-cards-and-prompt-formats/meta-llama-3.
ToolkenGPT  We implement ToolkenGPT with Llama 3 and train it on the same annotated datasets as we train Tecton, using the same sets of tools and the same hyperparameter combinations for direct comparability.
@@ Appendix D Out-of-distribution Datasets
Appendix E Prompts
In the reasoning phase, we prompt Tecton to generate an answer to a math reasoning question step by step, aided by in-context exemplars. Prompts D and D illustrate how we elicit tool choice in the meta-reasoning phase, for Tecton-score and Tecton-generate respectively. In both cases, we show the model in-context exemplars, followed by the current question and the current partial answer. The latter consists of the lines of text generated during the previous iterations of Tecton, if any. In Prompt D, the placeholder [FIXED EXEMPLARS] is replaced with six in-context exemplars extracted from the training set. The number of options in each exemplar matches the number of available options in the current sample. In Prompt D, we have [DYNAMIC EXEMPLARS] that are retrieved to demonstrate the tools currently in the candidate set. Here, the partial answer also includes the tokens generated during the current iteration, up to the sequence position where the first tool is found among the $k$ top probability tokens. The hints are prepended to the partial answer in this setting.

### v2

@@ Appendix C Baselines Implementation
All baselines are built upon Llama 3 8B Instruct.
Llama 3 8B Instruct.
We measure our base model’s performance in the zero-shot setting. We experiment with CoT zero-shot prompting (prepending to the question the instruction Let’s think step by step) but find that just using the raw question as input results in better performance.
Chain-of-Thought (CoT).
We extract from the training set six pairs of questions and answers demonstrating maths operations and their use. For comparability, we use the same exemplars as in the implementation of ToolkenGPT and Tecton. Note that Llama 3 8B Instruct achieves $79.6\%$ accuracy Dubey et al. (2024) on the non-enhanced version of GSM8K when prompted in few-shot CoT fashion. Our experiments show that it performs significantly worse ($37.3\%$) on GSM8K-XL, highlighting the difficulty of the enhanced dataset.
Trice.
Trice is a two-stage system where instruction tuning is followed by reinforcement learning from environmental feedback (RLEF), leveraging LoRA at both stages. We train Trice using the same hyperparameters as in the original implementation, with the exception of the batch size in the first phase of training, reduced to 128 with 8 gradient accumulation steps due to the memory limitations of our hardware. Since we train on a single dataset (GSM8K-XL), our implementation corresponds to the setup referred to as Trice-split in the original paper. Note that Qiao et al. (2024) implement Trice on top of Alpaca Taori et al. (2023), ChatGLM Zeng et al. (2024a) and Vicuna Zheng et al. (2023). For comparability with Tecton and the other baselines, here we use Llama 3 as the base model. We thus adjust Trice’s prompt templates and special tokens to be consistent with Llama 3’s model card [^1] https://www.llama.com/docs/model-cards-and-prompt-formats/meta-llama-3.
ToolkenGPT.
Like Tecton, ToolkenGPT augments the output matrix of an LLM with additional special tokens, each representing a tool. Only the additional tokens are tuned while the rest of the weights remain frozen.We implement ToolkenGPT with Llama 3 and train it on the same annotated datasets as we train Tecton, using the same sets of tools and the same hyperparameter combinations for direct comparability.
@@ Appendix E Prompts
In the reasoning phase, we prompt Tecton to generate an answer to a math reasoning question step by step, aided by in-context exemplars. Prompts E and E illustrate how we elicit tool choice in the meta-reasoning phase, for Tecton-score and Tecton-generate respectively. In both cases, we show the model in-context exemplars, followed by the current question and the current partial answer. The latter consists of the lines of text generated during the previous iterations of Tecton, if any. In Prompt E, the placeholder [FIXED EXEMPLARS] is replaced with six in-context exemplars extracted from the training set. The number of options in each exemplar matches the number of available options in the current sample. In Prompt E, we have [DYNAMIC EXEMPLARS] that are retrieved to demonstrate the tools currently in the candidate set. Here, the partial answer also includes the tokens generated during the current iteration, up to the sequence position where the first tool is found among the $k$ top probability tokens. The hints are prepended to the partial answer in this setting.
Appendix F Efficiency
We compare the efficiency of our method with the fine-tuned baselines—Trice and ToolkenGPT.
F.1 Comparison with Trice
Tecton is more efficient than Trice, as it only requires tuning additional embeddings in the output layer, with minimal backpropagation. In contrast, Trice performs two rounds of finetuning each time updating LoRA modules throughout the model.
F.2 Comparison with ToolkenGPT
We train Tecton in the same way as the ToolkenGPT baseline, by only updating the additional token embeddings in the output layer. At inference, Tecton requires generating additional text for each line in the problem. Note that for Tecton-score this overhead is negligible, as we generate only one additional token per line.